# Uncertainty-Aware Explanations Through Probabilistic Self-Explainable Neural Networks

### v1

@@ 6 Conclusions
In this paper we introduced Prob-PSENN, a novel Prototype-Based Self-Explainable Neural Network paradigm relying on a probabilistic generalization of the prototypes, which enhances the capabilities of the model in two key aspects. First, unlike previous approaches, our model is not restricted to a fixed number or set of prototypes. This enables Prob-PSENN to learn much more diverse prototypes, and to consider all of them when classifying an input, which represents a more flexible and robust approach. Second, our probabilistic approach equips Prob-PSENNs with tools to capture and harness the uncertainty in both the prediction and the explanation of the network. This novel feature makes it possible to detect when the model is providing uncertain outcomes, substantially increasing the reliability of the model. We experimentally demonstrate how Prob-PSENNs achieve more informative, robust and trustworthy explanations than their non-probabilistic counterparts, while keeping a competitive predictive accuracy.
@@ 7 Limitations and future work
As future work, inspired by recent advances in Bayesian Neural Networks Neal (2012); Gal (2016); Goan and Fookes (2020), we plan to develop a Bayesian formulation for the distribution over the prototypes, which might enable a more theoretically-grounded approach for uncertainty quantification. Furthermore, we also believe that a fully-probabilistic model (i.e., placing a probability distribution for the rest of the network parameters as well) could enhance the quantification of the uncertainty in Prob-PSENNs, or even induce new notions of uncertainty, such as the encoding or class-representation uncertainty. Apart from that, while Prob-PSENNs have demonstrated high predictive and explanatory performance, we found it challenging to model probability distributions over the prototypes in high-dimensional spaces, and, therefore, resort to relatively low dimensional latent spaces in our experiments. We plan to address this limitation in the future, in order to scale the proposed approach to higher-dimensional datasets.
@@ Acknowledgements
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 834115), the Spanish Ministry of Science, Innovation and Universities (project PID2022-137442NB-I00, grant FPU19/03231, and BCAM Severo Ochoa accreditation CEX2021-001142-S), and the Basque Government (projects KK-2023/00012 and IT1504-22).
@@ Appendix A Discussion on previous PSENN approaches
As introduced in Section 1, the development of PSENNs arises as an alternative to black-box DNNs and posthoc explanations, with the goal of ensuring a human-understandable decision process by construction, while keeping a competitive predictive performance. This is achieved by integrating a transparent case-based reasoning as the core element, so that the prediction of the model is determined based on the similarity between the input and a set of prototypical representations of the output classes (see Section 2 for further details).
@@
The way in which such prototypes are defined has led to different types of PSENNs in the literature. In the seminal work of Li et al. Li et al. [2018], mirrored by recent follow-up works Gautam et al. [2022], Joo et al. [2023], the prototypes represent complete characterizations of each class or concept (e.g., full representations of the digits in handwritten digit recognition). In contrast, other follow-up works aim to capture more abstract Alvarez-Melis and Jaakkola [2018] or specific features (e.g., prototypical parts or patches) Chen et al. [2019], Hase et al. [2019], Nauta et al. [2021] as prototypes. However, an important drawback of these approaches is that the prototypes are restricted to being samples or sample parts of (encoded) training inputs Alvarez-Melis and Jaakkola [2018], Chen et al. [2019], Hase et al. [2019], Nauta et al. [2021]. This implies a flexibility loss regarding the end-to-end optimization of the prototypes, as well as a transparency loss when surrogate prototypes (e.g., nearest training images) are selected to visualize the explanations Alvarez-Melis and Jaakkola [2018], Gautam et al. [2022]. For these reasons, we followed the same spirit as in Li et al. [2018], and thus employed the corresponding architecture as a baseline, in order to guarantee a fully transparent and end-to-end trainable architecture. We also remark that, while Gautam et al. [2022], Joo et al. [2023] also focus on learning a fixed set of complete, unconstrained, and decodable prototypes, the proposed architecture relies on a Variational Autoencoder Kingma and Welling [2014] as the autoencoding module. In contrast, as in Li et al. [2018], our approach does not require enforcing such a regularized latent space, hence representing a more general approach.
@@
Closer to our approach, concurrent works Joo et al. [2023], Wang et al. [2023] leverage the use of distributions over the prototypes to improve the consistency between image similarity and latent space location Joo et al. [2023] and to increase the representation power of the model Wang et al. [2023]. Contrary to our work, both approaches employ only the means of the learned distributions as prototypes, whereas our sampling-based inference process allows us to sample multiple sets of prototypes, avoiding the constraint of committing to a fixed set, and harnessing all of them for both the explanation and classification. Furthermore, the approach in Wang et al. [2023] follows the strategy of Alvarez-Melis and Jaakkola [2018], Chen et al. [2019], Hase et al. [2019] and rely on surrogate prototypes, replacing the distribution means with nearest training instances, with the corresponding flexibility and transparency loss discussed above. Another key difference between our approach and that in Joo et al. [2023] is that, in their work, the similarities between the input’s distributional embedding and the prototype distributions, which fed the final classification layer, is measured by considering the entire distributions, using distributional similarity metrics. However, as stated in Joo et al. [2023], such similarities do not form nor rely on observable prototypes, which is contrary to the purpose of explaining and justifying the classification. Instead, we advocate for a more transparent approach, in which the model predictions are solely derived from the distances to the sampled sets of prototypes, ensuring a complete and transparent justification of model outputs.
@@ Appendix B Implementation details
Our models have been implemented in Python, using TensorFlow Abadi et al. [2015] and TensorFlow Probability Dillon et al. [2017] packages. The experiments have been conducted on a cluster with Nvidia RTX A5000 GPUs and AMD EPYC 7252 8-Core CPUs. However, we point out that our models are amenable to being trained on commodity hardware. The architecture of the MLP Autoencoder module is composed of two dense layers with $1000$ hidden units and ReLU activation, followed by a dense layer with $l$ output neurons. A linear activation is used for the last layer of the encoder. The architecture of the CNN Autoencoder is composed of four convolutional layers, followed by a dense layer with $l$ output neurons and linear activation for the last layer of the encoder, in order to control the dimensionality of the latent space. The convolutional layers employ a $3\times 3$ kernel size, a stride value of 2, same zero padding. We use 32 channels for the first three layers and 10 channels for the fourth layer.
The experiments have been run for $\tau_{1}=1.0,\tau_{2}=10.0,\tau_{3}=0.05$, a batch size of 128 and 30 epochs. The Adam optimization method Kingma and Ba [2015] has been used to train the networks, using a learning rate of $0.001$. The final classification layer $W$ was set as $W=-I$, where $I$ represents the identity matrix of dimension $c$, thus ensuring that the association between the prototype $\bm{r}_{i}$ and the class $y_{i}$, $1\leq i\leq c$, is explicit and unequivocal. For all the datasets, the standard train and test splits have been used, with sizes 60000 and 10000, respectively, and a subset of 6000 training samples has been extracted as a validation set. The results on the baseline model Li et al. [2018] have been reproduced using the implementation released by the authors.
@@ Appendix C Complementary results on the MNIST dataset
In order to complement the results reported in Section 5, Figure 9 shows, for the MNIST dataset, 15 randomly sampled sets of prototypes from a Prob-PSENN with $l=5$ (left) and $l=10$ (right). As it can be observed, the increase in the dimensionality of the latent space makes it possible to capture prototype distributions with a higher degree of complexity in their structure, as well as to obtain more detailed decodings.
Figure 9: 15 random sets of prototypes, obtained with a Prob-PSENN with $l=5$ (left) and $l=10$ (right).
@@ Appendix D Results on the Fashion-MNIST dataset
This section includes complementary results on the Fashion-MNIST dataset. First, Figure 10 shows randomly sampled prototypes for a Prob-PSENN with $l=5$ (left) and $l=10$ (right). These figures demonstrate that the probabilistic redefinition of the prototypes introduced in Prob-PSENNs enables the model to learn, even for relatively low-dimensional latent spaces, diverse but still prototypical representations of each class.
Figure 10: Learned prototypes for the Fashion-MNIST dataset, with $l=5$ (left) and $l=10$ (right).
Furthermore, a visual comparison between the explanations provided by a non-probabilistic PSENN and a Prob-PSENN is provided in Figures 11 and 12. For the sake of completeness, the former figure includes the explanations provided by a PSENN with $m=15$ (left) and $m=50$ (right) prototypes. In both cases, the original input is shown in the top-left corner, and the prototypes are colored according to the classes returned by the model when the prototypes are classified, in order to make it clearer which class each prototype is most associated with. Notice that this connection is explicit and unequivocal by construction in our Prob-PSENNs, being the class $y_{i}$ represented by the prototype $r_{i}$, $1\leq i\leq c$, resulting in more transparent decision processes and easier-to-interpret explanations. As can be observed in Figures 11 and 12, the prototypes learned by the PSENNs are, even for $m=50$ prototypes, considerably less meaningful and varied than the ones captured by Prob-PSENN. Furthermore, the uncertainty metrics introduced in Section 3.3 enable us to quantify a low aleatoric and epistemic explanatory uncertainty for this input ($0.004$ and $0.16$, respectively), which increases the trustworthiness and reliability of the outcomes of Prob-PSENN.
Figure 11: Explanations provided by a PSENN with $m=15$ (left) and $m=50$ (right) for the Fashion-MNIST dataset. Both models classify the input (shown in the top-left corner of each grid) as the class pullover with a probability of $0.99$. Note that the prototypes are sorted based on their distances to the input (displayed on top of each figure), in increasing order.
Figure 12: Explanation provided by a Prob-PSENN, for the Fashion-MNIST dataset. The top row shows the distance distribution to the four closest classes (left), and the variability in the output probabilities (right). The bottom part shows the prototypes corresponding to the four closest classes, and the corresponding distances to the input. Note that the prototypes are ordered based on their distance, in increasing order.
@@ Appendix E Results on the K-MNIST dataset
This section contains additional results on the 10-class K-MNIST dataset Clanuwat et al. [2018], which includes handwritten representations of Japanese letters. Random images from this dataset are shown in Figure 13, from which it can be noticed a high variance in the representation of each letter. Randomly sampled prototypes from Prob-PSENNs with $l=5$ and $l=10$ are displayed in Figure 14. As can be assessed in the figure, despite the complexity and high variability of this dataset, Prob-PSENNs are capable of capturing the most frequent representations of each class, ensuring a sufficiently diverse and meaningful characterization of the classes.
@@
As for the previous datasets, a comparison between the explanations of PSENNs and Prob-PSENN for the K-MNIST dataset is provided in Figures 15 and 16. Notice how, also due to the higher variability in the representations of each class, the prototypes learned by a PSENN with $m=15$ prototypes (Figure 15-left) are not sufficiently representative of the classes, and, consequently, we need to resort to a large number of prototypes ($m=50$, Figure 15-right). However, even for $m=50$, the prototypes learned by PSENN are considerably less varied and informative than the ones sampled by Prob-PSENN, which evidences the higher flexibility of our approach, as well as its increased transparency and explanatory capabilities.
@@
Figure 13: Examples of the K-MNIST dataset Clanuwat et al. [2018], as reported in the original repository https://github.com/rois-codh/kmnist (CC BY-SA 4.0 license). Each row corresponds to one class. The first column of the figure shows the modern representation of each letter, as a reference.
@@
Figure 14: Sampled prototypes for the K-MNIST dataset, using a Prob-PSENN with $l=5$ (left) and $l=10$ (right).
@@
Figure 15: Explanations provided by a PSENN with $m=15$ (left) and $m=50$ (right) for the K-MNIST dataset. Both models classify the input (shown in the top-left corner of each grid) as the class 8 with a probability of $0.99$. Note that the prototypes are sorted based on their distances to the input (displayed on top of each figure), in increasing order.
@@
Figure 16: Explanation provided by a Prob-PSENN, for the K-MNIST dataset. The top row shows the distance distribution to the four closest classes (left), and the variability in the output probabilities (right). The bottom part shows the prototypes corresponding to the four closest classes, and the corresponding distances to the input. Note that the prototypes are ordered based on their distance, in increasing order.
Appendix F Results on the E-MNIST dataset

### v2

@@ 5 Conclusions
In this paper we introduced Prob-PSENN, a novel Prototype-Based Self-Explainable Neural Network paradigm relying on a probabilistic generalization of the prototypes, which enhances the capabilities of the model in two key aspects. First, unlike previous approaches, our model is not restricted to a fixed number or set of prototypes. This enables Prob-PSENN to learn much more diverse prototypes, and to consider all of them when classifying an input, which represents a more flexible and robust approach. Second, our probabilistic approach equips Prob-PSENNs with tools to capture the uncertainty in both the prediction and the explanation of the network. This novel feature makes it possible to detect when the model is providing uncertain outcomes, substantially increasing the reliability of the model. We experimentally demonstrated how Prob-PSENNs provide more informative and trustworthy explanations than their non-probabilistic counterparts, while achieving superior or competitive predictive accuracies. Moreover, we show that the accuracy of Prob-PSENNs can be improved by leveraging their novel uncertainty-quantification capabilities, further enhancing their reliability.
@@ 6 Limitations and future work
As future work, inspired by recent advances in Bayesian Neural Networks [30, 12, 15], we plan to develop a Bayesian formulation for the distribution over the prototypes, which might enable a more theoretically-grounded approach for uncertainty quantification. Furthermore, we also believe that a fully-probabilistic model (i.e., placing a probability distribution on the rest of the network parameters as well) could enhance the quantification of the uncertainty in Prob-PSENNs, or even induce new notions of uncertainty, such as the encoding or class-representation uncertainty. Apart from that, while Prob-PSENNs have demonstrated high predictive and explanatory performance, we found it challenging to model probability distributions over the prototypes in very high-dimensional spaces, and, therefore, resort to relatively low dimensional latent spaces in our experiments. We plan to address this limitation in the future, in order to scale the proposed approach to higher-dimensional datasets.
@@ Acknowledgements
MK received partial support from ELSA: European Lighthouse on Secure and Safe AI project (Grant No. 101070617 under UK guarantee) and the ERC under the European Union’s Horizon 2020 research and innovation program (FUN2MODEL, Grant No. 834115). JV, RS and JAL acknowledge support from the Spanish Ministry of Science, Innovation and Universities (Project PID2022-137442NB-I00 and BCAM Severo Ochoa Accreditation CEX2021-001142-S), and the Basque Government (Grant Nos. KK-2024/00030 and IT1504-22).
@@ Appendix A Discussion on previous PSENN approaches
As introduced in Section 1, the development of PSENNs arises as an alternative to black-box DNNs and posthoc explanations, with the goal of ensuring a human-understandable decision process by construction, while maintaining competitive predictive performance. This is achieved by integrating a transparent case-based reasoning as the core element, so that the prediction of the model is determined based on the similarity between the input and a set of prototypical representations of the output classes (see Section 2 for further details).
@@
The way in which such prototypes are defined has led to different types of PSENNs in the literature. In the seminal work of Li et al. [27], mirrored by recent follow-up works [13, 20], the prototypes represent complete characterizations of each class or concept (e.g., full representations of the digits in handwritten digit recognition). In contrast, other follow-up works aim to capture more abstract [3] or specific features (e.g., prototypical parts or patches) [7, 17, 29] as prototypes. However, an important drawback of these approaches is that the prototypes are restricted to being samples or sample parts of (encoded) training inputs [3, 7, 17, 29]. This implies a flexibility loss regarding the end-to-end optimization of the prototypes, as well as a transparency loss when surrogate prototypes (e.g., nearest training images) are selected to visualize the explanations [3, 13]. For these reasons, in order to guarantee a fully transparent and end-to-end trainable architecture, we followed the same spirit as in [27, 13], and used the corresponding approaches as references to build our probabilistic framework.
@@
Closer to our approach, concurrent works [20, 43] leverage the use of distributions over the prototypes to improve the consistency between image similarity and latent space location [20] and to increase the representation power of the model [43]. Contrary to our work, both approaches employ only the means of the learned distributions as prototypes, whereas our sampling-based inference process allows us to sample multiple sets of prototypes, avoiding the constraint of committing to a fixed set, and harnessing all of them for both the explanation and classification.
Furthermore, the approach in [43] follows the strategy of [3, 7, 17] and relies on surrogate prototypes, replacing the distribution means with nearest training instances, with the corresponding flexibility and transparency loss discussed above. Another key difference between our approach and that in [20] is that, in their work, the similarities between the input’s distributional embedding and the prototype distributions, which fed the final classification layer, are measured by considering the entire distributions, using distributional similarity metrics. However, as stated in [20], such similarities neither form nor rely on observable prototypes, which is contrary to the purpose of explaining and justifying the classification. Instead, we advocate for a more transparent approach, in which the model predictions are solely derived from the distances to the sampled sets of prototypes, ensuring a complete and transparent justification of model outputs.
@@ Appendix B Implementation details
The Prob-PSENN autoencoder module has been implemented as a CNN-based Variational Autoencoder, consisting of four convolutional layers with $4\times 4$ kernels, batch normalization, and LeakyReLU activations, followed by a $1\times 1$ convolutional layer that controls the dimensionality of the latent space. The decoder mirrors this architecture by means of deconvolutional layers. Further details on the architecture and the corresponding hyperparameters are provided in Figure B.12. The inputs values have been normalized to the range $[-1,1]$, and, for grayscale datasets, the input shapes have been upscaled from $28\times 28$ to $32\times 32$ using bilinear interpolation. For the SVHN dataset, random brightness adjustments and zoom transformations have been incorporated during training to improve generalization. The experiments have been run for $\tau_{1}=1.0,\tau_{2}=10.0,\tau_{3}=0.05,\tau_{4}=0.05$ and a batch size of 128. The models have been trained for 100 epochs on grayscale datasets and 30 epochs on the SVHN dataset. The Adam optimization method [22] has been used to train the networks, using a learning rate of $0.001$. The final classification layer $W$ has been set as $W=-I$, where $I$ represents the identity matrix of dimension $c$, thus ensuring that the association between the prototype $\bm{r}_{i}$ and the class $y_{i}$, $1\leq i\leq c$, is explicit and unequivocal. In order to ensure higher stability during the training process, the value of $\mathscr{L}_{INT}$ has been cut to the range $[-1,10^{3}]$. For all datasets, standard training and test partitions have been used, with 10% of the training data reserved for validation. The results of [27, 13] have been reproduced using the implementations released by the authors. Our code is publicly available at https://github.com/vadel/Prob-PSENN.
(a) MNIST, F-MNIST, K-MNIST
(b) SVHN
Figure B.12: VAE architectures used in Prob-PSENNs for (a) MNIST, F-MNIST, K-MNIST, and (b) SVHN.
@@ Appendix C Complementary results on the MNIST dataset
In order to complement the results reported in Section 4, Figure C.13 shows, for the MNIST dataset, 15 randomly sampled sets of prototypes from a Prob-PSENN with $l=5$ (left) and $l=10$ (right). As can be observed, the increase in the dimensionality of the latent space makes it possible to capture prototype distributions with a higher degree of complexity in their structure, as well as to obtain more detailed decodings.
Figure C.13: 15 random sets of prototypes, obtained with a Prob-PSENN with $l=5$ (left) and $l=40$ (right).
@@ Appendix D Results on the Fashion-MNIST dataset
This section includes complementary results on the Fashion-MNIST dataset. First, Figure D.14 shows randomly sampled prototypes for a Prob-PSENN with $l=5$ (left) and $l=40$ (right). These figures demonstrate that the probabilistic redefinition of the prototypes introduced in Prob-PSENNs enables the model to learn, even for relatively low-dimensional latent spaces, diverse and prototypical representations of each class.
Figure D.14: Sampled prototypes for the Fashion-MNIST dataset, using a Prob-PSENN with $l=5$ (left) and $l=40$ (right).
Furthermore, a visual comparison between the explanations provided by a non-probabilistic PSENN and a Prob-PSENN is provided in Figures D.15 and D.16. For the sake of completeness, the former figure includes the explanations provided by a PSENN with $m=15$ (left) and $m=50$ (right) prototypes. In both cases, the original input is shown in the top-left corner, and the prototypes are colored according to the classes returned by the model when the prototypes are classified, in order to make it clearer which class each prototype is most associated with. Notice that this connection is explicit and unequivocal by construction in our Prob-PSENNs, being the class $y_{i}$ represented by the prototype $r_{i}$, $1\leq i\leq c$, resulting in more transparent decision processes and easier-to-interpret explanations. As can be observed in Figures D.15 and D.16, the prototypes learned by the PSENNs are, even for $m=50$ prototypes, considerably less meaningful and varied than the ones captured by Prob-PSENN. Furthermore, the uncertainty metrics introduced in Section 3.3 enable us to quantify a low aleatoric and epistemic explanatory uncertainty for this input ($0.004$ and $0.16$, respectively), which increases the trustworthiness and reliability of the outcomes of Prob-PSENN.
Figure D.15: Explanations provided by a PSENN [27] with $m=15$ (left) and $m=50$ (right) for the Fashion-MNIST dataset. Both models classify the input (shown in the top-left corner of each grid) as the class pullover with a probability of $0.99$. Note that the prototypes are sorted based on their distances to the input (displayed on top of each figure), in increasing order.
Figure D.16: Explanation provided by a Prob-PSENN, for the Fashion-MNIST dataset. The top row shows the distance distribution corresponding to the four most likely classes (left), and the variability in the output probabilities (right). The bottom part displays the prototypes for the four most likely classes, along with their corresponding distances to the input. Note that the prototypes are ordered by their distance, in ascending order.
@@ Appendix E Results on the K-MNIST dataset
This section contains additional results on the 10-class K-MNIST dataset [8], which includes handwritten representations of Japanese letters. Random images from this dataset are shown in Figure E.17, from which it can be noticed a high variance in the representation of each letter. Randomly sampled prototypes from Prob-PSENNs with $l=5$ and $l=10$ are displayed in Figure E.18. As can be assessed in the figure, despite the complexity and high variability of this dataset, Prob-PSENNs are capable of capturing the most frequent representations of each class, ensuring a sufficiently diverse and meaningful characterization of the classes.
@@
As for the previous datasets, a comparison between the explanations of PSENNs and Prob-PSENN for the K-MNIST dataset is provided in Figures E.19 and E.20. Notice how, also due to the higher variability in the representations of each class, the prototypes learned by a PSENN with $m=15$ prototypes (Figure E.19-left) are not sufficiently representative of the classes, and, consequently, we need to resort to a large number of prototypes ($m=50$, Figure E.19-right). However, even for $m=50$, the prototypes learned by PSENN are considerably less varied and informative than the ones sampled by Prob-PSENN, which evidences the higher flexibility of our approach, as well as its increased transparency and explanatory capabilities.
@@
Figure E.17: Examples of the K-MNIST dataset [8], as reported in the original repository https://github.com/rois-codh/kmnist (CC BY-SA 4.0 license). Each row corresponds to one class. The first column of the figure shows the modern representation of each letter, as a reference.
@@
Figure E.18: Sampled prototypes for the K-MNIST dataset, using a Prob-PSENN with $l=5$ (left) and $l=40$ (right).
@@
Figure E.19: Explanations provided by a PSENN [27] with $m=15$ (left) and $m=50$ (right) for the K-MNIST dataset. Both models classify the input (shown in the top-left corner of each grid) as the class 8 with a probability of $0.99$. Note that the prototypes are sorted based on their distances to the input (displayed on top of each figure), in increasing order.
@@
Figure E.20: Explanation provided by a Prob-PSENN, for the K-MNIST dataset. The top row shows the distance distribution corresponding to the four most likely classes (left), and the variability in the output probabilities (right). The bottom part displays the prototypes for the four most likely classes, along with their corresponding distances to the input. Note that the prototypes are ordered by their distance, in ascending order.
Appendix F Ablation studies: complementary results
This Appendix complements the results from Section 4.4 by comparing the performance of Prob-PSENN with its deterministic counterpart, using the same architecture and training pipeline. The point-wise prototypes have been optimized using the loss function proposed in [27]. The results obtained with the deterministic PSENN are summarized in Table F.1, computed considering $l=20$ and $l=40$ dimensions for the latent space and various numbers of prototypes $m$. The results for Prob-PSENN are included in the last rows of the table (for $l=20$ and $l=40$), as reference. As can be observed, the performance of both approaches is nearly identical, with differences in average accuracy below 0.2%. These results further demonstrate that the advantages of our probabilistic model do not imply a reduction in its predictive performance, while bringing as benefits enhanced trustworthiness, uncertainty quantification and explanatory capabilities.
Table F.1: Predictive accuracy percentages (mean and standard deviation for 5 runs) for deterministic PSENNs using the same architecture and training pipeline as Prob-PSENNs. The best results for each approach (PSENN and Prob-PSENN) are highlighted in bold, in order to facilitate comparison and emphasize their similar performance.
Model Configuration MNIST F-MNIST K-MNIST
PSENN ($m\!=\!10$, $l\!=\!20$) 99.39 $\pm$ 0.04 92.11 $\pm$ 0.10 96.71 $\pm$ 0.12
PSENN ($m\!=\!10$, $l\!=\!40$) 99.42 $\pm$ 0.02 91.94 $\pm$ 0.25 96.79 $\pm$ 0.08
PSENN ($m\!=\!30$, $l\!=\!20$) 99.33 $\pm$ 0.10 90.92 $\pm$ 0.14 96.42 $\pm$ 0.18
PSENN ($m\!=\!30$, $l\!=\!40$) 99.29 $\pm$ 0.08 90.30 $\pm$ 0.51 96.72 $\pm$ 0.11
Prob-PSENN ($m\!=\!10$, $l\!=\!20$) 99.42 $\pm$ 0.05 91.91 $\pm$ 0.14 96.72 $\pm$ 0.21
Prob-PSENN ($m\!=\!10$, $l\!=\!40$) 99.38 $\pm$ 0.02 91.98 $\pm$ 0.24 96.61 $\pm$ 0.17
Next, in Table F.2 we report the results obtained with Prob-PSENNs for which VAEs have been replaced with vanilla AutoEncoders, matching all the remaining architectural details and the training pipeline. The results show that vanilla AutoEncoders lead to a performance drop ranging from approximately 0.1% to 1.2%, depending on the dataset. In addition, Figure F.21 visualizes different sets of randomly sampled prototypes from Prob-PSENNs with different backbone architectures, highlighting the advantage of regularized latent spaces for improving prototype quality in high-dimensional settings.
Table F.2: Comparison between the accuracy of Prob-PSENNs using vanilla AutoEncoder (AE) and Variational AutoEncoder (VAE) backbones.
Model Configuration MNIST F-MNIST K-MNIST SVHN
Prob-PSENN (AE, $l\!=\!10$) 99.26 $\pm$ 0.06 91.21 $\pm$ 0.48 95.58 $\pm$ 0.36 90.91 $\pm$ 0.26
Prob-PSENN (AE, $l\!=\!20$) 99.36 $\pm$ 0.05 91.76 $\pm$ 0.21 96.01 $\pm$ 0.12 90.72 $\pm$ 0.41
Prob-PSENN (AE, $l\!=\!40$) 99.35 $\pm$ 0.09 91.65 $\pm$ 0.30 96.00 $\pm$ 0.15 90.37 $\pm$ 0.22
Prob-PSENN (VAE, $l\!=\!20$) 99.42 $\pm$ 0.05 91.91 $\pm$ 0.14 96.72 $\pm$ 0.21 91.62 $\pm$ 0.56
Prob-PSENN (VAE, $l\!=\!40$) 99.38 $\pm$ 0.02 91.98 $\pm$ 0.24 96.61 $\pm$ 0.17 92.14 $\pm$ 0.24
(a) Prob-PSENN with a VAE backbone, $l=\mathbf{40}$.
(b) Prob-PSENN with an AE backbone, $l=\mathbf{10}$.
(c) Prob-PSENN with an AE backbone, $l=\mathbf{40}$.
Figure F.21: Visualization of randomly sampled prototype sets from Prob-PSENNs with the following backbone configurations: (a) VAE and $l=40$, (b) AE and $l=10$, and (c) AE and $l=40$. Prob-PSENNs with AE backbones provide a good performance in lower dimensions (e.g., $l=10$), but struggle to learn high-quality prototype distributions as dimensionality increases, whereas VAEs enable effective scaling to higher-dimensional spaces.
Appendix G Results on the E-MNIST dataset